# SINDI: an Efficient Index for Approximate Maximum Inner Product Search on Sparse Vectors

### v1

@@ 5 $\mathcal{D^{\prime}}\leftarrow\mathcal{D^{\prime}}\,\cup\vec{x}_{i}\_\alpha$
Algorithm 3 outlines the process for constructing the approximate version of the Sindi index. Given a sparse vector dataset $\mathcal{D}$, maximum dimension $d$, window size $\lambda$, and pruning ratio $\alpha$, the algorithm begins by initializing an empty set $\mathcal{D}^{\prime}$ to store pruned vectors (Line 1). For each vector $x_{i}$ in $\mathcal{D}$ (Line 2), its $\alpha$-mass subvector is derived and assigned to $\vec{x}_{i}^{\prime}\_\alpha$ (Line 3). The remaining steps are identical to the full-precision index construction process, with the exception that the approximate index stores the original dataset to enable reordering during retrieval.
Figure 6(a) shows that retaining less than half of the non-zero entries in a query can reduce the inner product error to nearly zero. This significantly reduces the search space, shrinking it by more than half. It indicates that the term lists corresponding to a few high-value non-zero entries in the query already cover most recall points. Therefore, Sindi applies Mass Vector Pruning to queries as well.
@@ 4.2. Reordering
Retaining a small portion of non-zero entries can preserve most of the inner product but fails to maintain the partial order of the full inner product. Using pruned results directly for recall reduces accuracy. However, experiments show that with enough candidates, the nearest neighbors are likely to be included. Figure 6(b) shows Recall 10@500 and Recall 10@10 under different pruning ratios for documents and queries. Retaining 16% of document entries and 30% of query entries achieves Recall 10@500=0.98, but Recall 10@10 is only 0.63. This suggests a two-step strategy: first, perform coarse recall with the pruned index to retrieve many candidates, then refine them using full inner product reordering for efficient AMIPS.
Input: Query $\vec{q}$, an inverted index $I$, query prune ratio $\beta$, reorder number $\gamma$, and $k$
Output: At most $k$ points in $\mathcal{D}$
@@
2HH is an empty minimumheapminimum\,heap;
3 $R$ is an empty minimumheapminimum\,heap;
@@
8while !$H.empty()$ do
9 $i,dis\leftarrow H.pop()$;
10 $dis^{\prime}\leftarrow\delta(\vec{x}_{i},\vec{q})$;
11 if $dis^{\prime}>R.min()$ or $R.len()<k$ then
@@ 10 $dis^{\prime}\leftarrow\delta(\vec{x}_{i},\vec{q})$;
Dataset $\|\mathcal{D}\|$ $avg\,\|x_{i}\|$ nqnq $avg\,\|q\|$ $d$ Sparsity Size (GB) avglavg\,l Model Language
@@
Detail of Algoruthm 4 Algorithm 4 details the query procedure for the approximate Sindi index. Given a query $q$, an inverted index $I$, query pruning ratio $\beta$, re-ranking threshold $\gamma$, and the number of nearest neighbors $k$, the algorithm starts by initializing two empty min-heaps, $H$ and $R$, for storing coarse recall candidates and final top-$k$ neighbors, respectively (Lines 1-2). The $\beta$ ratio subvector $q^{\prime}$ of the query $q$ is then generated (Line 3), and Algorithm 2 is invoked to retrieve $\gamma$ coarse recall results into $H$ (Line 4). While $H$ is not empty (Line 5), the algorithm processes its top element by extracting the stored vector ID $i$ and partial inner product score disdis (Line 6). The full inner product between $x_{i}$ and the original query $q$ is calculated as $dis^{\prime}$ (Line 7). The tuple $(i,\,dis^{\prime})$ is inserted into heap $R$, maintaining its size at most $k$ by removing the smallest element if necessary (Lines 8-11). This process continues until all coarse recall candidates in $H$ are processed.
5. Experimental Study
5.1. Experimental Settings
Datasets Table 3 presents the datasets used in this experimental evaluation. The experiments cover real-world datasets with varying languages and training models, ranging in size from 1M to 10M vectors. The English datasets were trained using the Splade model, while the Chinese datasets were trained using the BGE-M3 model, which was developed internally by Ant Group. Notably, the Chinese datasets have significantly higher dimensionality compared to the English datasets, primarily due to the larger vocabulary size in Chinese. Table 3 also includes the average number of non-zero entries per vector ($avg\,\|x_{i}\|$) and the average number of vectors per inverted list (avglavg\,l), which can serve as references for selecting specific pruning parameters. Additionally, the experiments incorporate randomly generated datasets, in which both the number of non-zero entries and the corresponding values follow a random distribution. Table 3 further provides the sparsity of each dataset, which quantifies how sparse the data is. The sparsity of a dataset $\mathcal{D}$ is calculated as: sparsity = $1-\frac{\sum_{x\in\mathcal{D}}\|x\|}{\|\mathcal{D}\|\cdot d}$, where $\|x\|$ denotes the number of non-zero entries in vector $\vec{x}$ (as defined in Definition 2.1), dd is the maximum dimension, and $\|\mathcal{D}\|$ is the total number of vectors in $\mathcal{D}$.
  • Seismic (seismic, ): A sparse vector index based on inverted lists.
  • BMP (bmp, ; sp, ): A dynamic pruning strategy for learning sparse vector retrieval. It divides the original dataset into fine-grained blocks and generates a maximum value vector for each block to evaluate whether the block should be queried.
  • HNSW (hnsw, ): A graph-based index designed for dense vector search. We adjusted the data format and distance computation to adapt it for sparse vectors.
  • PyAnns: The open-source champion of the BigANN Benchmark 2023 Sparse Track. It is built on HNSW and incorporates quantization, query pruning, and rerank strategies.
  • SOSIA (sosia, ): A sparse vector index based on min-hash.
@@ 5.1. Experimental Settings
Parameter Settings We used the optimal parameters for each algorithm to ensure a fair evaluation. The parameter selection either follows the recommendations from the original authors or is determined through grid search.
Performance Metrics We evaluate the index construction time, index size, recall, and QPS for all baselines. Since approximate methods require a trade-off between query efficiency and accuracy, we use throughput to measure query efficiency, which is defined as the number of queries processed within a specific time interval. For a given query $\vec{q}$, the result of Approximate Maximum Inner Product Search (AMPIS) is denoted as $R=\{\vec{x}_{1},\vec{x}_{2},\dots,\vec{x}_{k}\}$. Let $R^{*}=\{\vec{x}^{\ast}_{1},\vec{x}^{\ast}_{2},\dots,\vec{x}^{\ast}_{k}\}$ denote the exact top-$k$ results obtained via Maximum Inner Product Search (MIPS). The recall is computed as follows: $Recall=\frac{|R\cap R^{*}|}{|R^{*}|}$. We specifically evaluate the recall metrics Recall@50 and Recall@100.
The experiments are conducted on a server with an Intel(R) Xeon(R) Platinum 8269CY CPU @ 2.50GHz and 512GB memory. We implement Sindi in C++, and compile it with g++ 10.2.1, -Ofast flag, and AVX-512 instructions enabled.
@@ 5.2.1. Recall and QPS
Figure 8 analyzes the relationship between recall (Recall@50 and Recall@100) and the QPS performance of various algorithms on a single-threaded setup. For each algorithm, we report its best performance across all tested parameter configurations.
[…]
These results demonstrate that SINDI consistently achieves SOTA performance across datasets of various languages, models, and distributions.
@@ 5.2.2. Index Size and Construction Time
Figure 9 summarizes the index size and construction time for Sindi, Seismic, and PyAnns across three datasets. Sindi demonstrates the lowest construction cost across all datasets.
[…]
5.3. Parameters
5.3.1. The Impact of $\alpha$
This section explores how the document pruning parameter $\alpha$ affects Sindi’s performance. The parameter $\alpha$ determines the proportion of high-mass non-zero entries retained. For the MsMarco and NQ datasets, $\alpha$ is tested from 0.4 to 0.8, with a step size of 0.1. For the AntSparse dataset, $\alpha$ ranges from 0.7 to 1, with a step size of 0.05. Figure 10 presents the changes in recall and QPS as $\alpha$ increases under fixed $\beta$ and $\gamma$. On the MsMarco dataset, recall improves and QPS decreases as $\alpha$ grows, but both changes slow down when $\alpha$ becomes larger. On the AntSparse dataset, recall also improves slowly, but QPS drops more rapidly. When $\alpha$ is small, increasing it retains more high-scoring non-zero entries, which boosts recall significantly. As $\alpha$ becomes larger, the additional retained entries have less impact, making recall improvements slower. For the AntSparse dataset, the scores of non-zero entries have smaller variance and are closer to zero compared to English datasets. This causes more non-zero entries to be retained as $\alpha$ increases, leading to a faster decrease in QPS compared to the MsMarco dataset.
5.3.2. The Impact of sparsity
We analyze the performance of Sindi under varying dataset sparsity levels. sparsity measures the proportion of non-zero entries in a dataset. For a fixed number of non-zero entries, larger dimensions result in higher sparsity. To investigate this, we generated five random datasets, each with 1M vectors and an average of 120 non-zero entries per vector. The dataset dimensions were set to [10,000; 30,000; 50,000; 70,000; 100,000], increasing sparsity progressively. Figure 11 compares the performance of Sindi and Seismic across these datasets on Recall@50 = 90% and Recall@50 = 99%.
@@ 5.3.2. The Impact of sparsity
The results show that as sparsity increases, both Sindi and Seismic achieve higher QPS at the same recall levels. This is because higher sparsity allows the IVF structure to partition data more effectively. As sparsity increases, each term list cotains fewer vectors, which reduces the number of vector candidates that need to be searched during a query. Consequently, the number of inner product computations decreases, leading to improved QPS. Crucially, since all nearest neighbors are still captured within the lists, recall is not affected. In contrast, graph-based indexes face challenges with higher sparsity. Sparse high-dimensional data leads to weaker connectivity between nodes in the proximity graph, resulting in degraded search performance.Thus, IVF provide a natural advantage for sparse datasets compared with graph structure.
@@ 5.4.1. The Impact of Pruning Method
Figure 12 illustrates the performance of different pruning strategies on the SPLADE-FULL and AntSparse datasets. The experiments evaluate all pruning strategies under the same $\beta$ and $\gamma$ settings, while varying $\alpha$ to measure Recall and QPS. The results demonstrate that Mass Ratio Pruning achieves the best performance, followed by Vector Number Pruning and List Pruning, with the lowest performance observed when combining List Pruning and Mass Ratio Pruning.
This is because Mass Ratio Pruning effectively preserves the non-zero entries that contribute the most to inner product computation, resulting in more true nearest neighbors being retained during the partial inner product stage. In contrast, List Pruning restricts the size of posting lists for each dimension, which creates two problems: some lists contain too few documents, retaining small-value entries, while others have too many documents, removing large-value entries. As a result, List Pruning is unsuitable for Sindi. In comparison, Seismic employs List Pruning since it computes the full inner product for all vectors within the lists, thereby avoiding significant losses in accuracy. However, when both List Pruning and Mass Ratio Pruning are applied simultaneously, more non-zero entries are discarded, leading to further decreases in recall.
@@ 5.4.2. The Impact of Reorder
To investigate the impact of the reordering strategy on the performance of Sindi, we compared the differences between using and not using reordering. Experiments were conducted on the SPLADE-FULL and AntSparse datasets. We set the query cut ratio $\beta$ to 0.2 and reordering number $\gamma$ top 500. The non-reordering strategy did not require storing dataset.
The results are shown in Figure 13. We evaluated the query time and Recall@50 for both strategies under various $\alpha$. For the reordering strategy, the query time includes accumulation time and reordering time, while the non-reordering strategy only includes accumulation time. Since the parameter $\gamma$ is fixed, the reorder time remains relatively constant. However, as $\alpha$ increases, the accumulation time grows accordingly. Although reorder time accounts for only a small portion of the overall query time, it significantly improves recall. For example, on the SPLADE-FULL dataset, when $\alpha=0.6$, the accumulation time is 17099 ms, and the reorder time is 3553 ms. Despite this, the recall improves substantially from 0.71 to 0.97. This demonstrates the substantial efficiency improvement achieved by the reordering strategy.
The reordering strategy demonstrates its effectiveness for two reasons. First, it focuses on computing only a small subset of non-zero entries that contribute the most to the inner product, significantly reducing computations. Second, the partial inner product derived from high-mass entries can largely preserve the true ranking order of the inner product. This ensures that a small subset of candidate vectors is sufficient to include the true nearest neighbors, improving both efficiency and accuracy.

### v2

@@ 5 $\mathcal{D^{\prime}}\leftarrow\mathcal{D^{\prime}}\,\cup\vec{x}^{\prime}_{i}$
Algorithm 3 outlines the construction of the approximate Sindi index. Given a sparse dataset $\mathcal{D}$ of maximum dimension $d$, window size $\lambda$, and pruning ratio $\alpha$, the algorithm first initializes an empty set $\mathcal{D}^{\prime}$ to store pruned vectors (Line 1). For each vector $\vec{x}_{i}\in\mathcal{D}$ (Line 2), its $\alpha$-mass subvector $\alpha\text{-}mass(\vec{x}_{i})$ is computed and assigned to $\vec{x}^{\prime}_{i}$ (Line 3), which is then added to $\mathcal{D}^{\prime}$. The remaining steps invoke PreciseSindiConstruction on $\mathcal{D}^{\prime}$ (Line 5), followed by returning both the inverted index $I$ and the original dataset $\mathcal{D}$ (Line 6) so that reordering can be performed during retrieval.
Figure 6(a) shows that retaining under half of a query’s non-zero entries reduces the inner product error to nearly zero, cutting the search space by more than half. This suggests that posting lists from a few high–value non-zero entries in a query already cover most of the recall. Therefore, Sindi applies Mass Ratio Pruning to queries: given $\beta\in(0,1]$, the pruned query is denoted $\beta\text{-}mass(\vec{q})$ and is used in coarse retrieval.
@@ 4.2. Reordering
Retaining only a small portion of non-zero entries preserves most of the inner product but may disrupt the partial order of full inner products. Using such pruned results directly for recall degrades accuracy. Nevertheless, experiments show that with enough candidates, the true nearest neighbors are often included. Figure 6(b) reports Recall 10@500 and Recall 10@10 under different pruning ratios for documents and queries. Retaining 20% of document entries and 15% of query entries yields Recall 10@500 = 0.98 but Recall 10@10 = 0.63. This motivates a two-step strategy: (1) perform coarse recall with the pruned index to retrieve $\gamma$ candidates into a min-heap $H$; (2) compute the full inner products for all candidates in $H$ to refine the final top-$k$ results for efficient AMIPS. The second stage is reordering.
Input: Query $\vec{q}$, inverted index $I$, query prune ratio $\beta$, reorder number $\gamma$, and $k$
Output: Top-$k$ points in $\mathcal{D}$ (at most $k$)
@@
[…]
3 $R\leftarrow$ empty max-heap
@@
8
9while $H\neq\emptyset$ do
10 $i,dis\leftarrow H.pop()$
11 $dis^{\prime}\leftarrow 1-\delta(\vec{x}_{i},\vec{q})$
12 if $dis^{\prime}<R.max()$ or $R.len()<k$ then
@@ 11 $dis^{\prime}\leftarrow 1-\delta(\vec{x}_{i},\vec{q})$
Algorithm 4 shows the search procedure of the approximate Sindi index. Given a query $\vec{q}$, inverted index $I$, pruning ratio $\beta$, reordering size $\gamma$, and target $k$, the algorithm initializes an empty min-heap $H$ for coarse candidates and a max-heap $R$ for the final top-$k$ results (Lines 1–2). It first derives the $\beta$-mass subvector $\vec{q}^{\prime}$ (Line 3) and invokes PreciseSindiSearch to obtain $\gamma$ coarse candidates stored in $H$ (Line 4). While $H$ is not empty (Line 5), the best coarse candidate (i,dis)(i,\,dis) is popped (Line 6), its exact distance to $\vec{q}$ is computed as $dis^{\prime}$ (Line 7), and $(i,\,dis^{\prime})$ is inserted into $R$ if it improves the current set or if $R$ contains fewer than $k$ elements (Line 8). If $R$ exceeds size $k$, its worst element is removed (Lines 9–10). After all candidates are processed, $R$ is returned as the final result (Line 12).
5. Experimental Study
5.1. Experimental Settings
Datasets. Table 3 summarizes the datasets used in our experiments, covering: (i) English datasets from the MsMarco (Nguyen et al., 2016) passage ranking benchmark (including SPLADE-1M and SPLADE-FULL) and the NQ (Kwiatkowski et al., 2019) (Natural Questions) benchmark, all trained with the Splade model; (ii) Chinese dataset AntSparse, real business data from Ant Group trained with the BGE-M3 (Chen et al., 2024) model, which has higher dimensionality due to the larger Chinese vocabulary; (iii) Random datasets with non-zero entry dimensions and values drawn from a uniform distribution. For each dataset, Table 3 reports the average number of non-zero entries per vector ($\mathrm{avg}~\|\vec{x}_{i}\|$), average vectors per inverted list ($\mathrm{avg}~l$), and sparsity. The sparsity of $\mathcal{D}$ is: $\textit{sparsity}=1-\frac{\sum_{\vec{x}\in\mathcal{D}}\|\vec{x}\|}{\|\mathcal{D}\|\cdot d}$.
  • Seismic (Bruch et al., 2024): A sparse vector index based on inverted lists.
  • SOSIA (Zhao et al., 2024): A sparse vector index using min-hash signatures.
  • BMP (Mallia et al., 2024; Carlson et al., 2025): A dynamic pruning strategy for learning sparse vector retrieval. It divides the original dataset into fine-grained blocks and generates a maximum value vector for each block to evaluate whether the block should be queried.
  • HNSW (Malkov and Yashunin, 2020): A graph-based index originally for dense vectors; we modify the data format and distance computation to support sparse vectors.
  • PyANNS (PyANNS Contributors, 2025): The open-source champion of the BigANN Benchmark 2023 Sparse Track. It is built on HNSW and incorporates quantization, query pruning, and rerank strategies.
Parameter Settings. We use the optimal parameters for each algorithm to ensure a fair comparison. Parameter choices either follow the recommendations from the original authors or are determined via grid search. Details are shown in Table 4.
@@ 5.1. Experimental Settings
Dataset $\|\mathcal{D}\|$ $\mathrm{avg}~\|x_{i}\|$ nqnq $\mathrm{avg}~\|q\|$ $d$ Sparsity Size (GB) $\mathrm{avg}~l$ Model Language
@@
Table 4. Construction Parameter Settings
Algorithm Construction Parameter Settings
Sindi SPLADE-FULL: $\alpha\in$ {0.2, 0.3, 0.4, $\mathbf{0.5}$, 0.6, 0.7, 0.8}
AntSparse_10M: $\alpha\in$ {0.7, 0.75, 0.8, $\mathbf{0.85}$, 0.9, 0.95, 1}
Seismic SPLADE-FULL: $\lambda=6000$, $\beta=0.067$, $\alpha=0.4$
AntSparse_10M: $\lambda\in$ {5000, 6000, 7000, 8000, 9000, $\mathbf{10000},50000\}$, $\beta=0.1$, $\alpha=0.5$
PyANNS max_degree = 32, ef_construction = 1000
Sosia $m=150$, $l=40$
Bmp $l=16$
Hnsw max_degree = 32, ef_construction = 1000
Performance Metrics. We evaluate index construction time, index size, recall, and queries per second (QPS) for all baselines. For a query $\vec{q}$, let $R=\{\vec{x}_{1},\dots,\vec{x}_{k}\}$ denote the AMIPS results, and $R^{\ast}=\{\vec{x}^{\ast}_{1},\dots,\vec{x}^{\ast}_{k}\}$ denote the exact MIPS results. Recall is computed as $\mathrm{Recall}=\frac{\|R\,\cap\,R^{\ast}\|}{\|R^{\ast}\|}$. We specifically report Recall@50 and Recall@100. Since approximate methods trade off efficiency and accuracy, we also report throughput, defined as the number of queries processed per second.
Environment. Experiments are conducted on a server with an Intel Xeon Platinum 8269CY CPU @ 2.50GHz and 512 GB memory. We implement Sindi in C++, compiled with g++ 10.2.1 using -Ofast and AVX-512 instructions.
@@ 5.2.1. Recall and QPS
Figure 8 shows the relationship between recall (Recall@50 and Recall@100) and single-threaded QPS for all algorithms. For each method, we report the best results across all tested parameter configurations.
On both English and Chinese datasets, Sindi achieves the highest QPS at the same recall levels. When Recall@50 is 99%, on SPLADE-1M, the QPS of Sindi is $2.0\times$ that of Seismic and $26.4\times$ that of PyANNS; on SPLADE-FULL, it is $4.16\times$ and $10.0\times$ higher, respectively. When Recall@100 is 98% on SPLADE-FULL, Sindi attains $1.9\times$ the QPS of Seismic and $3.2\times$ that of PyANNS.
On the Chinese AntSparse-10M dataset encoded by the BGE-M3 (Chen et al., 2024) model, Sindi also performs best. Fixing Recall@50 at 97%, its QPS is $2.5\times$ that of Seismic, the recall of PyANNS is limited due to the high sparsity of the dataset.
On RANDOM-5M, generated uniformly at random, the uniform distribution yields very few common non-zero dimensions between vectors, causing graph-based methods (PyANNS, HNSW) to suffer severe connectivity loss and low recall. The clustering effectiveness of Seismic is also sensitive to data distributions, resulting in marked degradation. In contrast, Sindi is unaffected by data distribution and achieves the best performance, with QPS exceeding Seismic by an order of magnitude.
Overall, these results demonstrate that Sindi consistently delivers state-of-the-art performance across datasets with diverse languages, models, and distributions.
@@ 5.2.2. Index Size and Construction Time
Figure 9 summarizes the index size and construction time of Sindi, Seismic, and PyANNS on the two largest datasets (SPLADE-FULL and AntSparse-10M), showing that Sindi has the lowest construction cost.
Seismic, which stores summary vectors for each block, yields the largest index size; on AntSparse-10M, its size is $3.8\times$ that of Sindi. By contrast, the graph index construction of PyANNS requires numerous distance computations to find neighbors, resulting in a construction time $71.5\times$ that of Sindi on SPLADE-FULL. In comparison, Sindi mainly sorts each vector’s non-zero entries for pruning, keeping the cost low and enabling rapid index building.
@@ 5.3.1. The Impact of $\alpha$
This experiment examines how the document pruning parameter $\alpha$, which controls the proportion of high-mass non-zero entries retained, affects Sindi’s performance. A larger $\alpha$ retains more non-zero entries per vector, potentially increasing recall but also raising the search cost. We vary $\alpha$ from 0.4 to 0.8 (step 0.1) on SPLADE-FULL, and from 0.7 to 1.0 (step 0.05) on AntSparse, keeping $\beta$ and $\gamma$ fixed.
Figure 10 shows that, on MsMarco, recall rises and QPS drops as $\alpha$ increases, with both trends flattening at higher $\alpha$. In the lower $\alpha$ range, recall improves rapidly with moderate QPS loss, while in the higher range, recall gains slow and QPS stabilizes. On AntSparse, recall also grows more slowly at large $\alpha$, but QPS declines more steeply.
The slower recall gain at high $\alpha$ is due to the saturation effect in Section 4.1: once enough non-zero entries are kept, further additions barely reduce inner product error. The sharper QPS drop on AntSparse comes from its lower variance of non-zero values, which leads to more retained entries for the same $\alpha$ increase, and thus more postings to scan.
5.3.2. The Impact of sparsity
We evaluate the performance of Sindi under varying dataset sparsity levels. For a fixed $\mathrm{avg}\,\|\vec{x}\|$, larger $d$ yields higher sparsity. To examine its impact, we generate five synthetic datasets ($\|\mathcal{D}\|=1\text{M}$, $\mathrm{avg}\,\|\vec{x}\|=120$) with $d\in\{10\mathrm{k},\,30\mathrm{k},\,50\mathrm{k},\,70\mathrm{k},\,100\mathrm{k}\}$, thereby increasing sparsity gradually. Each dataset is generated uniformly at random, where both the positions and the values of non-zero entries follow a uniform distribution. Figure 11 compares Sindi and Seismic at Recall@50 = 90% and Recall@50 = 99%.
() Recall@50=90%
(a) Recall@50=99%
As sparsity increases, both Sindi and Seismic achieve higher QPS at the same recall because the IVF index produces shorter inverted lists (average length $\mathrm{avg}\,l$ decreases), reducing the number of candidate non-zero entries $\lVert q\rVert\,l$ to be visited. Since all true nearest neighbors still reside in the probed lists, recall remains unaffected.
@@ 5.4.1. The Impact of Pruning Method
Figure 12 illustrates the performance of different pruning strategies on the SPLADE-FULL and AntSparse datasets. All strategies are evaluated under the same $\beta$ and $\gamma$ settings, while varying $\alpha$ to measure Recall and QPS. Mass Ratio Pruning achieves the best overall performance, followed by Vector Number Pruning and List Pruning, with the lowest performance observed when combining List Pruning and Mass Ratio Pruning.
The advantage of Mass Ratio Pruning lies in its ability to preserve the non-zero entries that contribute most to the inner product, thereby retaining more true nearest neighbors during the coarse recall stage. In contrast, List Pruning limits the posting list size for each dimension, which can result in two issues: some lists become too short and keep mainly small-value entries, while others remain too long and remove large-value entries. As a result, List Pruning is less suitable for Sindi. Seismic, however, uses List Pruning because it computes the full inner product for all vectors in the lists, avoiding large accuracy losses. When List Pruning is combined with Mass Ratio Pruning, even more non-zero entries are discarded, further reducing recall.
@@ 5.4.2. The Impact of Reorder
() SPLADE-FULL
(a) AntSparse-10M
To investigate the impact of the reordering strategy on the performance of Sindi, we compared the cases with and without reordering on the SPLADE-FULL and AntSparse datasets. The query prune ratio $\beta$ was set to 0.2, and the reordering number $\gamma$ was fixed at 500. For the document prune ratio $\alpha$, we varied it from 0.3 to 0.6 in steps of 0.1 on SPLADE-FULL, and from 0.7 to 1.0 in steps of 0.1 on AntSparse-10M. The non-reordering strategy does not require storing the dataset. We evaluated both strategies using query time and Recall@50 as performance metrics.
The results are shown in Figure 13. For the reordering strategy, the query time includes both accumulation time and reordering time, whereas the non-reordering strategy includes only accumulation time. Since $\gamma$ is fixed, the reorder time remains relatively constant. As $\alpha$ increases, the accumulation time also increases accordingly. Although reorder time accounts for only a small portion of the total query time, it yields a substantial recall improvement. For example, on SPLADE-FULL with $\alpha=0.6$, the accumulation time is 17099 ms and the reorder time is 3553 ms ($\approx 17.2\%$ of the total), yet recall improves from 0.71 to 0.97. This demonstrates the clear benefit of incorporating the reordering strategy.
The reordering strategy is effective for two main reasons. First, it focuses computation on a limited set of non-zero entries that contribute most to the inner product, greatly reducing unnecessary operations. Second, the partial inner products derived from high-value entries largely preserve the true ranking order, ensuring that small $\gamma$ is sufficient to contain the true nearest neighbors, thereby improving both efficiency and accuracy.